Calculate ROA and ROE ratios of your selected company for 3 years based on DuPont Ratio. Analyze the trend and explain the reason of the increase/decrease in those ratios.

Provide a table of three years Actual and five years forecast of income Statements of your selected company. Include the assumptions that you used for your forecast. For example how did you calculate future revenues, expenses, and other items and why?

Breakdown 2019 2018 2017 2016
Total Revenue 4'509'000 4'313'200 4'088'400 3'900'900
Cost of Revenue 1'517'600 1'434'000 1'425'300 1'376'200
Gross Profit 2'991'400 2'879'200 2'663'100 2'524'700
Operating Expenses
Selling General and Administrative - - - 0
Total Operating Expenses 2'398'400 2'242'500 2'065'700 1'959'500
Operating Income or Loss 593'000 636'700 597'400 565'200
Interest Expense 35'000 18'800 15'600 14'100
Income Before Tax 561'100 620'600 582'800 554'100
Income Tax Expense 49'200 133'500 130'300 134'300
Income from Continuing Operations 511'900 487'100 452'500 419'800
Net Income 509'600 485'100 450'700 419'700
Net Income available to common shareholders 509'600 485'100 450'700 419'700
EBITDA 918'900 818'900 767'400 719'700
Intangible Assets 603,700 602,100 610,000 620,800
Total non-current assets 5,065,100 4,316,800 4,194,500 4,054,300
Total Assets 8,040,800 7,249,800 6,975,600 6,428,800
Liabilities and stockholders' equity
Liabilities
Current Liabilities
Current Debt 505,500 12,300 9,200 316,400
Accounts Payable 233,900 214,200 217,300 180,400
Total Current Liabilities 1,689,700 1,028,100 1,049,800 1,258,900
Non-current liabilities
Long Term Debt 498,500 998,700 998,200 748,900
Deferred taxes liabilities 541,500 467,000 444,200 433,500
Other long-term liabilities 5,900 6,600 7,700 7,600
Total non-current liabilities 1,680,900 1,735,300 1,730,800 1,495,900
Total Liabilities 3,370,600 2,763,400 2,780,600 2,754,800
Stockholders' Equity
Common Stock 24,300 24,300 24,100 23,700
Retained Earnings 4,982,200 4,655,400 4,246,200 3,743,800
Total stockholders' equity 4,659,800 4,477,300 4,186,300 3,667,200
Total liabilities and stockholders' equity 8,040,800 7,249,800 6,975,600 6,428,800
Answer #1
Return on Assets (ROA)
ROA=(Net Income /Sales)*(Sales/Average Total Assets)
ROA=Net Profit Margin*Asset Turnover
Average Total Assets=(Year Beginning Asset +Year Ending asset)/2
Average Total Assets 2017 6702200 (6428800+6975600)/2
Average Total Assets 2018 7112700 (6975600+7249800)/2
Average Total Assets 2019 7645300 (7249800+8040800)/2
2017 2018 2019
A Net Income 450700 485100 509600
B Sales 4088400 4313200 4509000
C Average Total Asset 6702200 7112700 7645300
D=A/B Net Profit Margin 0.110239 0.112469 0.11301841
E=B/C Asset Turnover 0.610009 0.606408 0.58977411
ROA=D*E Return on Assets 0.067247 0.068202 0.06665533
Return on Assets Percentage 6.72% 6.82% 6.67%
Return on Equity (ROE)
ROE=(Net Income /Sales)*(Sales/Average Total Assets)*(AverageTotal Assets/Average Shareholder Equity)
ROE=ROA*(AverageTotal Assets/Average Shareholder Equity)
ROE=ROA*Financial Leverage
Average Shareholder Equity 2017 3926750 (3667200+4186300)/2
Average Shareholder Equity 2018 4331800 (4186300+4477300)/2
Average Shareholder Equity 2019 4568550 (4477300+4659800)/2
2017 2018 2019
F ROA 0.067247 0.068202 0.06665533
G Average Total Assets 6702200 7112700 7645300
H Average Shareholder Equity 3926750 4331800 4568550
I=G/H Financial Leverage 1.706806 1.641973 1.67346313
ROE=F*I Return on Equity 0.114777 0.111986 0.11154524
Return on Equity Percentage 11.48% 11.20% 11.15%